Cantel Company produces cleaning compounds for both commercial and household customers. Some of these products are produced as part of a joint manufacturing process. For example, GR37, a coarse cleaning powder meant for commercial sale, costs $1.90 a pound to make and sells for $2.00 per pound. A portion of the annual production of GR37 is retained for further processing in a separate department where it is combined with several other ingredients to form SilPol, which is sold as a silver polish, at $5.00 per unit. The additional processing requires 1/5 pound of GR37 per unit; additional processing costs amount to $4.10 per unit of SilPol produced. Variable selling costs for SilPol average $0.30 per unit. If production of SilPol were discontinued, $2,000 of costs in the processing department would be avoided. Cantel has, at this point, unlimited demand for, but limited capacity to produce, product GR37.

Required 1. Calculate the minimum number of units of SilPol that would have to be sold in order to justify further processing of GR37. 2. Assume that the cost data reported for GR37 are obtained at a level of output equal to 6,600 pounds, which is the maximum that the company can produce at this time. What is the expected operating income (loss) under each of the following scenarios: (a) all available capacity is used to produce GR37, but no SilPol; (b) 6,000 units of SilPol are produced, with the balance of capacity devoted to the production and sale of GR37; (c) 10,000 units of SilPol are produced, with the balance of capacity devoted to the production and sale of GR37; and (d) 12,000 units of SilPol are produced, with the balance of capacity devoted to the production and sale of GR37.
Answer #1

Details provided in the question:

Cost of producing GR37 = $1.9 per pound

Selling price of GR37 = $2 per pound

Selling price of SilPol = $5 per unit

Quantity of GR37 required for producing per unit of GR37 = 1/5 pound of GR37

Additional processing cost = $4.10 per unit

Variable selling cost = $0.30 per unit

Amount that can be avoided by discontinuing SilPol = $2000

Solution:

1. Calculation of minimum number of SilPol to be sold in order to justify further production:

In order to justify further production the company will have to recover the additional cost of $2000 which it incurs solely due to production of SilPol.

Total variable cost of producing 1 unit of SilPol = Cost of GR37 + Additional Processing cost + Variable selling cost

= [(1/5)*1.9] + 4.10 + 0.30

= $ 4.78

Contribution per unit of SilPol = Selling Price - Total Variable cost

= 5 - 4.78

= $0.22

Minimum number of SilPol to be sold in order to justify further production = Additional Firedcost Contribution per unitof Sil Pol

= \frac{2000}{0.22}

= ~9091 units

Therefore, Minimum number of SilPol to be sold in order to justify further production is 9091 units.

2. (a) All available capacity is used to produce GR37

Total pounds of GR37 that will be produced will be = 6600 pounds

Profit per unit of GR37 = Selling price - Cost = $2 - $1.9 = $0.10

Total profit on sale of 6600 pounds of GR37 = $0.10 * 6600 = $660

2(b) 6000 units of SilPol and balance capacity devoted to GR37

Total quantity of GR37 required to produce 6000 units of SilPol = 6000*(1/5) = 1200 pounds

Balance quantity of GR37 that can be sold = 6600 - 1200 = 5400 pounds

Profit / (loss) on sale of 6000 units of SilPol = Contribution - Fixed Cost

= ($ 0.22 * 6000 units) - $2000

= 1320 - 2000

= (680)

Loss on sale of 6000 units of SilPol = $680

Profit on sale of 5400 pounds of GR37 = 5400 pounds * $0.10 = $540

Total profit / (loss) = (680) + 540 = (140)

2(c) 10000 units of SilPol and balance capacity devoted to GR37

Total quantity of GR37 required to produce 10000 units of SilPol = 10000*(1/5) = 2000 pounds

Balance quantity of GR37 that can be sold = 6600 - 2000 = 4600 pounds

Profit / (loss) on sale of 10000 units of SilPol = Contribution - Fixed Cost

= ($ 0.22 * 10000 units) - $2000

= 2200 - 2000

= 200

Profit on sale of 10000 units of SilPol = $200

Profit on sale of 4600 pounds of GR37 = 4600 pounds * $0.10 = $460

Total profit / (loss) = 200 + 460 = 660

2(d) 12000 units of SilPol and balance capacity devoted to GR37

Total quantity of GR37 required to produce 12000 units of SilPol = 12000*(1/5) = 2400 pounds

Balance quantity of GR37 that can be sold = 6600 - 2400 = 4200 pounds

Profit / (loss) on sale of 12000 units of SilPol = Contribution - Fixed Cost

= ($ 0.22 * 12000 units) - $2000

= 2640 - 2000

= 640

Profit on sale of 12000 units of SilPol = $640

Profit on sale of 4200 pounds of GR37 = 4200 pounds * $0.10 = $420

Total profit / (loss) = 640 + 420 = 1060